# Neural Stereo Video Compression with Hybrid Disparity Compensation

### v1

@@ 3.2.2 HDC for Cross-View Entropy Modeling
In this progressive coding scheme, each new slice’s entropy modeling is conditioned on previously encoded information. Specifically, for slice $\bm{y}_{n}^{M}$ (where $M\in\{L,R\}$, denotes the current view and $\overline{M}$ denotes the opposite view), the entropy estimation network $Est_{n}^{M}$ integrates three types of priors to predict the probability distribution of the quantized slice $\bm{\hat{y}}_{n}^{M}$. First, intra-view priors, which include previously encoded slices from the same view, $\bm{\hat{y}}_{1\sim n-1}^{M}=\{{\bm{\hat{y}}_{1},\dots,\bm{\hat{y}}_{n-1}}\}^{M}$. These capture the already-known information from earlier slices of the current view. Second, cross-view priors, which consist of encoded slices from the opposite view, $\bm{\hat{y}}_{1\sim q}^{\overline{M}}=\{\bm{\hat{y}}_{1},\dots,\bm{\hat{y}}_{q}\}^{\overline{M}}$, where $q=n-1$ for $M=L$ (left view) and $q=n$ for $M=R$ (right view). In other words, when encoding a left-view slice, we use all slices up to index $n-1$ from the right view as cross-view priors, and for a right-view slice we use up to index $n$ from the left view. The HDC mechanism is applied to these cross-view priors to align them with the current view, compensating for disparity via disparity-aware shifting and cross-attention fusion. Third, spatial–temporal priors, which capture spatial and temporal context for the current slice, are denoted as $\bm{\Phi}_{1\sim n}^{M}$. These priors are obtained by fusing the hyperprior features with multi-scale context features ($\bm{Ctx}_{1\sim 3}^{M}$), as illustrated in Fig. 3 (c). The spatial–temporal priors provide additional cues from the broader image context and temporal reference frames. By incorporating these three sets of priors, $Est_{n}^{M}$ can more accurately estimate the probability distribution for $\bm{\hat{y}}_{n}^{M}$, leading to improved entropy modeling for that slice.
The entropy estimation process within $Est_{n}^{M}$ is illustrated in Fig. 3 (b). First, the HDC mechanism performs cross-view alignment: it takes the available encoded slices from the opposite view $\bm{\hat{y}}_{1\sim q}^{\overline{M}}$ and aligns them to the current view $M$ using disparity compensation. This alignment uses spatial–temporal prior of the current view as an anchor to guide the disparity-aware shifting and cross-attention, following Sec.3.2.1. However, unlike the earlier HDC-FER module, this alignment step omits any downsampling or upsampling operations and directly outputs the aligned cross-view priors (denoted as $\bm{s}_{1\sim q}^{\overline{M}}=\{\bm{s}_{1},\dots,\bm{s}_{q}\}^{\overline{M}}$).
@@
[…]
@@
Figure 4: Rate-distortion curves. The results are evaluated on the KITTI 2012, KITTI 2015 and Nagoya [3] datasets in terms of Bpp-PSNR.
3.3 Objective Function
@@
4 Experiments
4.1 Evaluation Datasets.
@@ 4 Experiments
Cityscapes. The Cityscapes dataset [16] comprises 2975, 500, and 1525 stereo sequences for training, validation, and testing, respectively. Each sequence contains 30 frames at a resolution of 2048$\times$1024. Following [11, 24], we crop frames to 1920$\times$704 by removing the top, left, and bottom regions. We use the test set of Cityscapes to analyze performance, with the Group of Pictures (GOP) set to 30, matching the sequence length.
KITTI. The KITTI 2012 [19] and KITTI 2015 [46] datasets contain 195 and 200 stereo sequences, respectively, with each sequence comprising 21 frames. Following [11, 24], we crop frames to 1216$\times$320 for testing. The GOP is set to 21 for evaluation.
Nagoya. Unlike Cityscapes and KITTI, which focus on autonomous driving, the Nagoya dataset [3] provides general-scene multi-view video sequences. To assess the generalization capability of our model, we select two standard MV-HEVC benchmark sequences [53], Kendo and Balloons, from the Nagoya MPEG-FTV project [3]. We evaluate our method on the first 96 frames, using the first and third views from both sequences. The GOP is set to 32.
@@
Ours -53.13 -55.97 -54.72 -27.96 -19.92
Table 1: BD-rate (%) comparison in RGB colorspace measured with PSNR. The anchor is MV-HEVC.
4.2 Experiment Protocols
Evaluation Metrics. We evaluate image reconstruction quality using two widely adopted metrics: Peak Signal-to-Noise Ratio (PSNR) and Multi-Scale Structural Similarity Index Measure (MS-SSIM) [58]. Additionally, we measure compression efficiency using bits per pixel (Bpp). During compression, the first frame of each GOP (I-frame) is encoded using the learned image compression model proposed by He et al. [23], while subsequent frames (P-frames) are compressed using our method.
Implementation Details. We train four models, each corresponding to a different rate-distortion trade-off, determined by the Lagrange multiplier $\lambda$ ($\lambda=256,512,1024,2048$). The training process includes pretraining a single-view variant of our model on the Vimeo-90K dataset [63] and then finetuning on the Cityscapes dataset [16] following previous methods [11, 25]. Further details of training procedures and hyperparameters are provided in Sec. A in the appendix.
Figure 5: (a) and (b) present the ablation studies of our proposed modules on the Cityscapes dataset [16]. (c) shows the comparison with previous SOTA NSVCs (LSVC [11] and LLSS [25]) under the same backbone on the KITTI 2012 dataset [19].
4.3 Comparison with State-of-the-art Methods
Baseline Methods. Following previous works [25, 11], we compare our model with various single-view and stereo video compression methods, including traditional codecs, neural-based video compression baselines, and state-of-the-art NSVC approaches. Specifically, we evaluate against traditional methods HEVC [52], MV-HEVC [53] and neural-based video compression methods FVC [27], as well as neural-based stereo video compression methods LSVC [11] and LLSS [25].
For traditional baselines, we use the HM-16.20 [1] implementation with the lowdelay_P_main preset for HEVC and the HTM-16.3 [2] implementation with the baseCfg_2view configuration for MV-HEVC. The results for HEVC, MV-HEVC, FVC, LSVC, and LLSS on the Cityscape [16] and KITTI [19, 46] datasets are obtained via email correspondence with the authors, ensuring consistency with previous comparisons.
Additionally, we compare with DCVC-TCM*, a computationally efficient variant of DCVC-TCM [50] and the backbone of our framework. DCVC-TCM* reduces complexity by using fewer channels and a lighter motion estimation network [54], while preserving its core architecture. Our framework extends DCVC-TCM* by integrating HDC-FER and HDC-EM, further improving compression efficiency. For fair evaluation, DCVC-TCM* follows the same training pipeline described in Sec. 4.2, with pretraining on Vimeo-90K [63] and fine-tuning on Cityscapes [16].
Rate-Distortion Performance. The rate-distortion (RD) curves in Fig. 4 compare our proposed method against baseline approaches in terms of Bpp-PSNR across four datasets: KITTI 2012 [19], KITTI 2015 [46], and the Kendo and Balloons video sequences from the Nagoya dataset [3]. As the figure illustrates, our model consistently achieves state-of-the-art performance across all datasets, demonstrating a significant improvement in rate-distortion efficiency over existing methods. LSVC and LLSS are not included in the comparison on Kendo and Balloons due to the unavailability of their pretrained models.
Tab. 1 provides a detailed quantitative analysis using BD-Rate [8] as the evaluation metric, calculated based on Bpp-PSNR with MV-HEVC [53] as the anchor. On the autonomous driving datasets (i.e., Cityscapes, KITTI 2012, and KITTI 2015), our method achieves an average BD-rate saving of approximately 54.61% compared to the anchor MV-HEVC, and around 26.81% compared to the previous SOTA NSVC method LLSS. We would also like to emphasize that on general-scene multi-view video sequences (i.e., Kendo and Balloons), our method achieves an average BD-rate saving of approximately 23.91% compared to the anchor MV-HEVC, and outperforms our single-view baseline DCVC-TCM* by about 24.54% on average. This demonstrates the superiority of exploiting disparity redundancy in indoor stereo scenes, whereas previous NSVC methods (e.g., LSVC and LLSS) typically only showcase such benefits in auto-driving scenarios The results indicate that our proposed approach outperforms traditional and neural-based stereo video codecs, achieving lower BD-Rate values across all four datasets. Performance comparisons using MS-SSIM as the distortion metric and a model complexity analysis of the proposed method against baseline methods are detailed in Sec. B and Sec. C of the appendix.
4.4 Ablation Study
Effectiveness of HDC-FER and HDC-EM. To evaluate the individual and combined effectiveness of our proposed HDC-FER and HDC-EM modules, we conduct an ablation study based on the single-view baseline model DCVC-TCM* using the Cityscapes dataset. Specifically, the variant DCVC-TCM* + HDC-FER integrates HDC-FER modules into the feature extraction and reconstruction process of the baseline method DCVC-TCM*, enabling enhanced feature representation during the encoding and decoding stages to simultaneously compress the two views of stereo frames. DCVC-TCM* + HDC-EM replaces the original single view entropy module of DCVC-TCM* with the proposed cross-view HDC-EM entropy model, which leverages information from both views to achieve more accurate entropy estimation for each view in an autoregressive manner. The results presented in Fig. 5 (a) demonstrate that integrating HDC-FER alone (DCVC-TCM* + HDC-FER) achieves a BD-rate reduction of 6.03%, and substituting the entropy module with HDC-EM (DCVC-TCM* + HDC-EM) results in a BD-rate reduction of 26.59%. When both modules are simultaneously integrated into the model (DCVC-TCM* + HDC-EM + HDC-FER), the total BD-rate reduction is further improved to 32.17%, thereby confirming the significant benefits of employing cross-view information and enhanced feature representation for compression performance.
Effectiveness on Implicit and Explicit Compensation in HDC. To validate the effectiveness of the implicit and explicit compensation within our HDC mechanism, we conduct an ablation study on the Cityscapes dataset, focusing on the “attention score” (i.e., implicit cross-attention) and the shift operation (i.e., explicit pixel displacement). First, removing the “attention score” from both HDC-EM and HDC-FER modules results in cross-view features being extracted solely from the shifted feature volumes, without the guidance of attention weighting. As shown in Fig. 5 (b), this modification leads to a 13.75% increase in BD-rate, highlighting the significant role of the implicit cross-attention mechanism in improving compression efficiency. Next, to assess the contribution of explicit pixel displacement, we disable the shift operation in both modules, thereby allowing feature interaction between views only through simple channel attention. This results in a 9.08% increase in BD-rate, as illustrated in Fig. 5 (b), demonstrating the effectiveness of explicit pixel displacement in facilitating cross-view feature extraction.
Effectiveness of Single-View NVC Backbone. To rigorously assess the effectiveness of the proposed HDC-FER and HDC-EM modules, we take the KITTI 2012 dataset as an example to perform a controlled comparison by integrating them into the same backbone architecture as prior SOTA NVC methods, including LSVC [11] and LLSS [25]. Concretely, we start from the official pretrained FVC [27] model trained on the Vimeo-90K dataset [63], incorporate our modules into its framework, and finetune the resulting model on Cityscapes [16], following the protocol detailed in Sec. 4.2.
As presented in Fig. 5 (c), our method achieves superior performance over prior SOTA approaches under identical backbone settings, thereby substantiating the contribution of the proposed modules to compression efficiency. Moreover, when employing DCVC-TCM* as the backbone—which corresponds to the final configuration of our model—further performance gains are observed, reaffirming the scalability and general applicability of the proposed HDC-FER and HDC-EM modules.
5 Conclusion
In this work, we propose a novel neural stereo video compression (NSVC) framework that integrates a hybrid disparity compensation (HDC) strategy to effectively reduce cross-view redundancy. By combining explicit pixel displacement for structural priors with implicit cross-attention mechanisms for flexible feature alignment, HDC enhances both disparity compensation accuracy and optimization efficiency. We further incorporate HDC into two key modules: HDC-FER for cross-view feature extraction and reconstruction, and HDC-EM for cross-view entropy modeling, enabling more effective compression. Extensive experiments on both autonomous driving and general multi-view datasets demonstrate that our framework outperforms previous neural and traditional stereo video compression methods, highlighting its superiority in both efficiency and reconstruction quality.
Appendix
Stage P Frames Network Loss LR Iterations
Stage 2 4 All (Add cross-view priors) $D_{rec}^{*}\&R_{all}^{*}$ 1e-5 3e5
Stage 3 4 HDC-FER modules $D_{rec}^{*}\&R_{all}^{*}$ 1e-5 2e4
Stage 4 4 The full model $D_{rec}^{*}\&R_{all}^{*}$ 1e-5 2e5
Table S2: Training strategy
In the appendix, we provide additional details and experimental results to further support our framework. Section A describes the training strategy and experimental settings for both training and testing. Section B presents an extended rate-distortion analysis using BD-rate measured with MS-SSIM. To evaluate computational efficiency, Section C compares the FLOPs and parameter counts of our method with baseline approaches. Finally, Section D details the network architecture, including the integration of HDC-FER and HDC-EM into the motion and context compression modules.
Appendix A Detailed Training and Testing Settings
Training Strategy.
Our training process consists of four stages. First, following DCVC-TCM [50], we pretrain a single-view variant of our model on the Vimeo-90K dataset [63], excluding all HDC-FER modules and cross-view priors in entropy modeling. We then progressively incorporate our proposed HDC-FER and HDC-EM modules into the model while fine-tuning it on the Cityscapes dataset [16] using the Adam optimizer [30] with a batch size of 4 and a learning rate of $1\times 10^{-5}$. Throughout training, we optimize the cascaded distortion loss [44] $D_{\text{rec}}^{*}$, and the rate loss, $R_{\text{all}}^{*}$ over four consecutive P-frames, preserving gradients across these frames. The full training strategy is outlined in Table S2.
@@ Appendix A Detailed Training and Testing Settings
In Stage 2, we introduce cross-view priors into the entropy model by integrating HDC-EM and fine-tune for 300k iterations. In Stage 3, HDC-FER modules are incorporated and pretrained for 20k iterations while keeping other components fixed. Finally, in Stage 4, the full model (with both HDC-EM and HDC-FER) undergoes an additional 200k iterations of fine-tuning.
[…]
Training Settings
Following prior works [11, 25], we optimize our model using MSE and MS-SSIM as distortion metrics. The loss function is defined as $MSE(\bm{x}_{t}^{M},\bm{\hat{x}}_{t}^{M})$ for MSE optimization and $1-MSSSIM(\bm{x}_{t}^{M},\bm{\hat{x}}_{t}^{M})$ for MS-SSIM optimization.
For I-frame compression, we employ the learned image compression model ELIC [23], fine-tuning it on the Cityscapes [16] training set for 200 epochs with a batch size of 8. The rate-distortion trade-off parameter $\lambda$ is set to $\{0.008,0.016,0.016,0.15\}$ for MSE-based training and $\{3,12,12,40\}$ for MS-SSIM-based training.
For P-frame training, we pretrain our model on Vimeo-90K [63] using randomly cropped into 256×256. In later training stages, Cityscapes frames are cropped to $256\times 384$. To enhance model generalization, we apply random horizontal and vertical flipping as data augmentation.
Testing Settings.
We follow the preprocessing procedures of prior works [25, 11] for evaluating on Cityscapes [16], KITTI 2012 [19], and KITTI 2015 [46]. For Cityscapes, we crop 64 pixels from the top, 256 pixels from the bottom, and 128 pixels from the left to remove rectification artifacts and the ego vehicle, resulting in a resolution of $1920\times 704$. For the KITTI 2012 and KITTI 2015 datasets, frames are cropped to a size of $1216\times 320$ by removing pixels from the top and left. Additionally, we exclude video sequences 000127 and 000182 from KITTI 2012, and 000026 and 000167 from KITTI 2015 to maintain consistency with prior studies. For the newly introduced dataset Nagoya [3], no cropping is applied, and frames are kept at the original 1024 $\times$ 768 resolution.
Appendix B BD-rate measured by bpp-MS-SSIM
Method Cityscapes KITTI 2012 KITTI 2015 Kendo Balloons
MV-HEVC [53] 0.0 0.0 0.0 0.0 0.0
HEVC [52] 34.37 14.23 16.96 238.97 1449.67
FVC [27] -30.96 -40.83 -35.34
DCVC-TCM* [33] -35.13 -48.86 -38.46 -42.02 26.05
LSVC [11] -46.88 -51.60 -47.20
LLSS [25] -58.10 -55.71 -51.16
Ours -58.95 -60.03 -51.19 -55.49 -8.84
Table S3: BD-rate (%) comparison in RGB colorspace measured with MS-SSIM. The anchor is MV-HEVC.
We evaluate performance using MS-SSIM as the distortion metric, calculating BD-rate [8] with MV-HEVC [53] as the anchor. The results, presented in Table S3, show that our method outperforms all baseline approaches, achieving the best performance in terms of Bpp-MS-SSIM.
Appendix C Model complexity
Method Params(M) FLOPs(G)
DCVC-TCM [50] 10.55 735.95
DCVC-TCM* 9.68 431.24
LSVC [11] 50.50 1760.0

### v2

@@ III-C2 HDC for Cross-View Entropy Modeling (HDC-EM)
Figure 3: (a). Overview of the Hybrid Disparity Compensation module for Entropy Modeling (HDC-EM). The input is the latent feature pair $\{\bm{y}^{L},\bm{y}^{R}\}$ produced by the encoder modules of the Motion and Context Compression components (see the middle part of Fig. 1). Each latent feature $\bm{y}^{L}$ (resp., $\bm{y}^{R}$) is evenly divided along the channel dimension into N slices $\bm{y}_{1\sim N}^{L}$ (resp., $\bm{y}_{1\sim N}^{R}$). Meanwhile, spatial–temporal context features $\bm{\varPhi}_{1\sim N}^{L}$ (resp., $\bm{\varPhi}_{1\sim N}^{R}$) are generated using the network in (c). To estimate the distribution of each quantized latent slice $\bm{\hat{y}}^{L}$ (resp., $\bm{\hat{y}}^{R}$), we use both the context features and the accumulated losslessly compressed slices from the same and opposite views as priors. These are passed to the entropy estimation module $Est_{n}^{M}$ in (b) for accurate entropy coding. (b). The entropy estimation network $Est_{n}^{M}$ predicts the probability distribution for the n-th slice $\bm{\hat{y}}_{n}^{M}$ at the current view $M$, using the accumulated losslessly compressed slices from the same view $M$, (i.e., $\bm{\hat{y}}_{1\sim n-1}^{M}$) and from the opposite view $\overline{M}$ (i.e., $\bm{\hat{y}}_{1\sim q}^{\overline{M}}$, s.t., q=n1q=n-1 for $M=L$ and $q=n$ for $M=R$), along with the accumulated context features $\bm{\varPhi}_{1\sim n}^{M}$. To align the cross-view slices $\bm{\hat{y}}_{1\sim q}^{\overline{M}}$ with the current view, we apply the HDC alignment strategy, which shares a similar architecture to that shown on the right side of Fig. 1. (c) The spatial–temporal prior fusion network generates the context $\bm{\Phi}^{M}$ by fusing the hyperprior, extracted from $\bm{y}^{M}$ via the hyper encoder-decoder ($h_{a}$ and $h_{s}$), with the multi-scale features $\bm{Ctx}^{M}_{1\sim 3}$ (see the left part of Fig. 1). The resulting context is then divided into N channel-wise slices, denoted as $\bm{\Phi}_{1\sim N}^{M}$, following the same slicing operation used for latent features. The context fusion module adopts a structure similar to the temporal context encoder in DCVC-TCM [50], while $h_{a}$, $h_{s}$}, and the prior fusion module consist of several convolutional layers.
@@
In this progressive coding scheme, each new slice’s entropy modeling is conditioned on previously encoded information. Specifically, for the input slice $\bm{y}_{n}^{M}$ (where $M\in\{L,R\}$, denotes the current view and $\overline{M}$ denotes the opposite view), the entropy estimation network $Est_{n}^{M}$ integrates three types of priors to predict the probability distribution of the quantized slice $\bm{\hat{y}}_{n}^{M}$. First, intra-view priors, which include previously encoded slices from the same view, $\bm{\hat{y}}_{1\sim n-1}^{M}=\{{\bm{\hat{y}}_{1},\dots,\bm{\hat{y}}_{n-1}}\}^{M}$. These capture the already-known information from earlier slices of the current view. Second, cross-view priors, which consist of encoded slices from the opposite view, $\bm{\hat{y}}_{1\sim q}^{\overline{M}}=\{\bm{\hat{y}}_{1},\dots,\bm{\hat{y}}_{q}\}^{\overline{M}}$, where q=n1q=n-1 for $M=L$ (i.e., the left view) and $q=n$ for $M=R$ (i.e., the right view). In other words, when encoding a left-view slice, we use all slices up to index n1n-1 from the right view as cross-view priors, and for a right-view slice we use up to index $n$ from the left view. The HDC mechanism is applied to these cross-view priors to align them with the current view, compensating for disparity via disparity-aware shifting and cross-attention fusion. Third, spatial-temporal priors, which capture spatial and temporal context for the current slice, are denoted as $\bm{\Phi}_{1\sim n}^{M}$. These priors are obtained by fusing the hyperprior features with multi-scale context features $\bm{Ctx}_{1\sim 3}^{M}$ [^3] For the motion compression network, only hyperprior features are used; context features are not involved., which are derived from a motion compensation module (see the left part of Fig. 1), as illustrated in Fig. 3 (c). The spatial–temporal priors provide additional cues from the broader image context and temporal reference frames. By incorporating these three sets of priors, $Est_{n}^{M}$ can more accurately estimate the probability distribution for $\bm{\hat{y}}_{n}^{M}$, leading to improved entropy modeling for that slice.
The entropy estimation process within $Est_{n}^{M}$ is illustrated in Fig. 3 (b). First, the HDC mechanism performs cross-view alignment: it takes the available encoded slices from the opposite view $\bm{\hat{y}}_{1\sim q}^{\overline{M}}$ and aligns them to the current view $M$ using disparity compensation. This alignment uses spatial–temporal prior of the current view as an anchor to guide the disparity-aware shifting and cross-attention, similar to Sec. III-C1. In contrast to the previous HDC-FER module, this alignment step simplifies the process by excluding any downsampling, upsampling, or additive operations, and directly yields the aligned cross-view priors (denoted as $\bm{s}_{1\sim q}^{\overline{M}}=\{\bm{s}_{1},\dots,\bm{s}_{q}\}^{\overline{M}}$).
Figure 4: Rate-distortion (RD) curves. The results are evaluated on the KITTI 2012, KITTI 2015 and Nagoya [21] datasets in terms of Bpp-PSNR.
@@
where $\lfloor\cdot\rceil$ denotes rounding to the nearest integer. The resulting quantized slice $\bm{\hat{y}}_{n}^{M}$ is stored and treated as a known prior for subsequent entropy estimation steps, enabling the encoding process to adapt progressively as more slices (and cross-view information) become available.
@@
III-D Objective Function
@@
IV Experiments
IV-A Evaluation Datasets.
@@ IV Experiments
[…]
Figure 5: Subjective quality comparison on KITTI 2012 dataset.
[…]
Nagoya. In contrast to Cityscapes and KITTI, which focus on autonomous driving scenarios, the Nagoya dataset [21] provides multi-view video sequences of general scenes. To assess the generalization capability of our model, we selected two standard MV-HEVC benchmark sequences, Kendo and Balloons, available through the Nagoya MPEG-FTV project [21]. Evaluation was performed on the first 96 frames, utilizing the first and third views from both sequences. The frames were kept at their original 1024 × 768 resolution for testing. The GOP was set to 32 for evaluation. We test on the RGB videos which are obtained from the original YUV420 format with BT.709 conversion.
IV-B Experiment Protocols
Figure 6: Ablation study on rate-distortion (RD) performance by progressively integrating the proposed HDC-EM and HDC-FER modules into the baseline DCVC-TCM*.
Evaluation Metrics. To assess the quality of frame reconstruction, we adopt RGB-PSNR as the distortion metric. Compression efficiency is evaluated using bits per pixel (Bpp). Furthermore, BD-rate [61] is employed to provide a comprehensive quantitative comparison. For the initial frames of both views (i.e., the I-frames), compression is performed independently using the learned image compression model proposed by He et al. [58], while the proposed method is applied to the remaining frames (i.e., the P-frames).
TABLE I: BD-RATE (%) COMPARISON IN RGB COLORSPACE MEASURED WITH PSNR. THE ANCHOR IS MV-HEVC.
@@ IV-B Experiment Protocols
Ours -53.13 -55.97 -54.72 -19.52 -10.61
Implementation Details. We train all models by minimizing a rate-distortion objective, weighted by a Lagrange multiplier $\lambda$, as formulated in Eq. 6. Following [3], we set the maximum disparity value $D$ to 192, corresponding to the maximum disparity reported in the KITTI datasets [19, 20]. All models are implemented in PyTorch and optimized with the Adam optimizer[62].
To ensure stable optimization, we divide the training process into four stages. In Stage 1, following the training process used in DCVC-TCM [50], a single-view variant of our model (i.e., DCVC-TCM*)—excluding all HDC-FER modules and cross-view priors in HDC-EM entropy model—is pretrained on the Vimeo-90K dataset [63] with a batch size of 4. Subsequently, the model is fine-tuned on the Cityscapes dataset [60] with a fixed learning rate of $1\times 10^{-5}$ during Stages 2 to 4.
Figure 7: Ablation study on RD performance by removing the attention score and shift operation from the proposed HDC mechanism.
@@
In Stage 2, cross-view priors are introduced into HDC-EM entropy model, and the network is fine-tuned for 300k iterations. In Stage 3, HDC-FER modules are added and trained for 20k iterations, with all other parameters frozen. Finally, in Stage 4, the complete model (including both HDC-EM and HDC-FER) undergoes an additional 200k iterations of fine-tuning.
All training is conducted on two NVIDIA RTX 4090 GPUs. During training, input frames are cropped to $256\times 256$ in Stage 1 and $256\times 384$ in the subsequent stages.
IV-C Comparison with State-of-the-art Methods
Baseline Methods. Following [3, 4], we also compare our method against traditional video codecs HEVC [23], a traditional MVC method MV-HEVC [2], a neural video codec FVC [27] and two state-of-the-art NSVC methods LSVC [4] and LLSS [3]. Specifically, we adopt HM-16.20 [64] and HTM-16.3 [65] as the official reference implementations of HEVC and MV-HEVC following [4, 3], respectively. For HEVC, we use the encoder_lowdelay_P_main configuration, and for MV-HEVC, we adopt the baseCfg_2view setting. The results of LSVC [4] and LLSS [3] are reported from their original papers.
In addition, we compare our approach with DCVC-TCM*, a lightweight variant of DCVC-TCM [50], which serves as the backbone of our framework. DCVC-TCM* reduces model complexity by employing fewer channels and a lightweight motion estimation module [66], while maintaining the overall architecture. These modifications significantly lower the computational cost and reduce encoding latency, making it more suitable for resource-constrained environments. As shown in Tab. II, DCVC-TCM* achieves substantial improvements in computational efficiency and encoding speed. Our framework extends DCVC-TCM* by incorporating HDC-FER and HDC-EM modules to further enhance compression efficiency. For fair comparison, DCVC-TCM* is trained following the similar pipeline as our method, with pretraining on Vimeo-90K [63] and fine-tuning on Cityscapes [60].
TABLE II: Model complexity on $512\times 512$ stereo videos, evaluated under the settings of [3]. DCVC-TCM* refers to a lightweight version of DCVC-TCM [50], used as our baseline.
Method Params (M) FLOPs (G) Enc Speed Dec Speed
DCVC-TCM [50] 10.55 735.95 14.29fps 17.85fps
DCVC-TCM* 9.68 431.24 18.87fps 18.52fps
LSVC [4] 50.50 1760.0 6.25fps 8.33fps